OpenStudy (anonymous):

y = x - 2 y = 4x + 1 Solving systems by substitution .

OpenStudy (anonymous):

just put the value of y from the first equation in second: x-2=4x+1 solve it for x put value of x in any original equation - you will have y. Let me know if there is any questions

OpenStudy (anonymous):

how do i solve for x ?

OpenStudy (anonymous):

x-2=4x+1 put all x on one side: 4x-x+1=-2 -1 from both sides: 3x=-3 divide both sides by (3): x=-1 now, you put value of x say in first equation: y=x-2=-1-2=-3 so, y=-3 your answer should be: x=-1, y=-3

OpenStudy (anonymous):

thank you so much !